# Particle migration in areas of constricted flow

### v1

@@ Abstract
Cardiovascular diseases are a leading cause of death globally. Among them, some are linked to stenosis, which is an abnormal narrowing of blood vessels, as well as other factors. Smart drug delivery systems based on micro- and nanoparticles are a promising method to offer non/minimal-invasive therapeutic mechanisms. Here we investigate the propensity of particles with different shapes and sizes to drift laterally (marginate) towards an occlusion area in a two-dimensional (2D) parallel plate laminar flow using the Lattice-Boltzmann method (LBM). To verify the outcomes on both sides of the stenosis, a probability of adhesion to the borders was calculated. Analysis was done on the impact of wall-shear stress on both sides of the stenosis. Our results show that non-circular particles migrate in larger amounts and earlier than circular ones.
I Introduction
@@
Regarding the characterization of the blood, it is worth noting that, despite its non-Newtonian nature, for many applications blood can be considered a Newtonian fluid, provided that the wall shear stress observed in the regions of study are not too low [25, 26]. In such cases, the rheological differences in blood behavior are negligible and the Newtonian assumption holds true. Since our main concern is analyzing particle-fluid interaction and not the properties of blood itself, and since we do not consider red blood cells in this study (which are major players in the shear-thinning behavior of the blood), the use of a Newtonian model is justified.
II Numerical method
To study the motion of a particle cluster in a 2D constricted channel, the Lattice-Boltzmann method (LBM)[27, 24, 28] is adopted. The dynamics of the Newtonian fluid flow with an external force can be described by a single relaxation time via the Bhatnagar-Gross-Krook (BKG) equation [29],
@@ II Numerical method
Here $\rho$ and $\bf{u}$ are the fluid density and velocity, respectively; $\omega_{i}$ represents the weighting parameter of the model, where $\omega_{0}=4/9$, $\omega_{1-4}=1/9$ and $\omega_{5-8}=1/36$; ${\bf{c}}_{i}$ is the discrete velocity of the model, with $c_{0}=(0,0)$, $c_{1-4}=c(\cos[(i-1)\pi/2],\sin[(i-1)\pi/2])$, $c_{5-8}=2c(\cos[(2i-1)\pi/4],\sin[(2i-1)\pi/4])$; $c=\Delta x/\Delta t$ is the lattice speed, which was assumed to be one since $\Delta x=\Delta t=1$; $\Delta x$ is the lattice spacing; and $c_{s}=c/\sqrt{3}$ is the sound speed of the model. The last term in Eq. (1) represents the source term, which takes into account the force density driving the fluid as well as the forces exerted by particles [31]. Based on this scheme, a body force density $\bf{F}$ is related to $F_{i}$ as,
@@
The local fluid density, $\rho$, and velocity, $\bf{u}$, at the lattice node are given by $\rho=\sum_{i}f_{i}$ and $\rho{\bf{u}}=\sum_{i}f_{i}{\bf{c_{i}}}+\Delta t{\bf{F}}/2$. $\bf{F}$ is the sum of the forces acting on the particles. The LBM for a single-phase flow recovers the Navier-Stokes equation in the limit of small Knudsen number for incompressible fluids, with the fluid kinematic viscosity being $\nu=c_{s}^{2}(\tau-0.5)$ (in lattice units). The kinematic viscosity cannot be less than zero; therefore, the limitation $\tau>0.5$ must be satisfied.
@@
where $m_{p}$ and $I_{p}$ denote the mass and moment of inertia of the $p$ particle, ${\bf{v}}_{p}$ and $\Omega_{p}$ are the translational and rotational velocities of each particle, $\bf{F_{D}}$ represents the drag force, $\bf{F_{L}}$ is the lift force, and $\bf{T}$ the particle torque. Note that the effect of forces like virtual mass, Basset history and Magnus effect can be considered negligible. Other forces (Brownian, gravity, etc) acting on the particles are not considered in this paper. Particles boundary are discretized in $M$ nodes whose velocities ${\bf{v}}_{b}$ are given by,
@@
with ${\bf{r}}_{b}$ and ${\bf{r}}_{p}$ being the boundary node and particle positions, respectively. Finally, the drag and lift forces and the torque are given as,
@@
$\bf{T}$ $=$ $C_{D}\sum_{b}\left(({\bf{r}}_{b}-{\bf{r}}_{p})\times({\bf{u}}_{b}-{\bf{v}}_{p})\right)$ (10)
where $C_{D}$ and $C_{L}$ denote the drag and lift coefficients, respectively, and ${\bf{u}}_{b}={\bf{u}}({\bf{r}}_{b})$ is the fluid velocity at the boundary nodes. For the lift equation (9) we considered the Saffman force [32] for the sake of simplicity, although more complex models could be used [33].
@@
Figure 1 shows the 2D rectangular domain ($H\times L=400\times 4000$ cells) considered (see the Appendix A for a grid convergence study). A Poiseuille flow with a maximum velocity $u_{max}$ at $H/2$ was considered at the inlet and zero-flux boundary condition for the outlet. The bounce-back scheme was implemented on channel walls to simulate the no-slip condition on the fluid, capturing the wall and inertial effects on particle-fluid motion, as discussed by Basagaoglu et al. (2008)[39]. The constriction zone or stenosis is simulated as a Gaussian function of height $h_{0}$ and variance $h_{0}^{2}$ centered at $L/2$.
@@
Particles had initially the same velocity as the flow at the release location. In this paper, we focus specifically on circular and rectangular rigid particles. $N$ particles were released randomly far enough from the constriction zone and at $H/4$ cells far from the lateral boundaries (gray area in Fig. 1). Care was put in making sure there is no overlapping between them. Rectangles were set initially at random orientations. For a circular particle with radius $R$ rotating around its center of mass, the angular moment of inertia is $I_{p}=m_{p}R^{2}/2$, while for a rectangular particle with sides $l_{1}$ and $l_{2}$, $l_{2}\geq l_{1}$, $I_{p}=m_{p}(l_{1}^{2}+l_{2}^{2})/12$. We consider elastic collisions between two particles, or a particle and a wall. For the collision between circular particles we suppose a hard sphere collision, whereas for the collision between two rectangles we follow Qin et al. (2020) [34]. For the last case, the collision model takes into account whether the collision involves two vertices, a vertex and an edge or two edges. In this paper, we do not consider collisions between circular and rectangular particles. To compare results between both types of particles, we define the equivalent radius for the rectangle particles as $R_{equiv}=l_{1}(1+\gamma^{2})^{1/2}/2$ where the aspect ratio is defined as $\gamma=l_{2}/l_{1}$. The fluid is initialized with a density $\rho=1$, velocity $u_{max}=0.04$, Reynolds number $Re=Hu_{max}/\nu$ and a relaxation time $\tau=0.62$ (all given in lattice units).
III Results
III.1 Flow trajectories of particles
@@
The LBM was used to simulate the migration trajectories of particles as a function of their size and shape. Simulation results in Fig. 2 unveiled two distinct shape dependent particle behaviors in a constricted channel: (i) both circular and rectangular particles migrate towards the lateral boundaries; however, the rectangles do so before the circles. Additionally, (ii) rectangular particles migrate to both sides of the stenosis and in greater quantity than circles. Note that both types of particles create a wake, in which others are caught and therefore experience less drag, resulting in slower motion and eventual clustering. This is less pronounced for rectangles as they displayed non-zero cumulative angular rotations favoring their dispersion, unlike the circular particles.
@@
Particle migration to the edges depends on their size and shape as their response to drag and lift forces is different. Particles are considered to have migrated to the edges if the distance from any particle node to the boundaries is less than 5 cells and they are in the intervals $[L/2-L/8,L/2]$ before and $[L/2,L/2+L/8]$ after the stenosis. Fig. 4 shows the ratio of migrated particles $N_{migr}/N$ for circles and rectangles as a function of $R_{equiv}$ before (a) and after (b) the stenosis. Note that for both occlusion percentages, $N_{migr}$ is larger for the rectangular particles than for the circular ones. Small circular particles hardly see their trajectory modified and are instead confined in the center-region of the channel. Before the stenosis, $N_{migr}$ increases with $R_{equiv}$ in all cases, while the number of circles reaching the boundaries after the stenosis decreases as their size increases. Large particles have larger residence times, Fig. 4(c), as they were less frequently trapped by the recirculating vortices. Residence time increases with the size of the occlusion zone and the recirculating vortices.
Once particles cross the stenosis, Fig. 4(b), circles are dragged by the central flow, where the velocity is at a maximum ($N_{migr}$ decreases with increasing $R_{equiv}$). For the rectangles, lift forces increase and they have a higher probability of migrating to the boundaries. Residence times do not change appreciably with $R_{equiv}$, but, in contrast to the previous case, decrease as the percentage of occlusion increases, Fig. 4(d). Very large particles can collapse/stop the flow at the stenosis by stacking [35].
@@
The influence of the occlusion percentage on particle migration is shown in Fig. 5. For both types of particles with similar values of $R_{equiv}$, the number of particles that migrate to the boundaries increases with an increasing occlusion. Wall-shear stress upstream of the stenosis increases with $h_{0}$, favoring migration of particles to the boundaries; conversely, post-stenosis wall-shear stress diminishes due to the presence of a recirculation zone, thus reducing the rate of particle migration, especially for circular particles, which are more prone to be dragged by the flow. The driving forces for margination are significantly altered by the constriction.
@@
Fig. 6 illustrates the influence of the number of released circular particles $N$ and the Reynolds number on the ratio of particles that migrated to the boundaries. Only results containing circular particles are displayed; earlier results demonstrating rectangular particles indicate that $N_{migr}$ is greater than that of circular particles. As expected, $N_{migr}$ increases with $N$ before and after the stenosis, but increasing $N$ leads to larger slow flow at the particles wake and clogging at the constriction, thus reducing the ratio of particles that migrate to the lateral boundaries with respect to the number of emitted particles. This is more evident after the stenosis. On the other hand, increasing the Reynolds number by increasing $u_{max}$ leads to a linear increase of $N_{migr}$, as the drag and lift forces are proportional to $Re$.
@@ III.2 Probability of adhesion
There are numerous models for calculating the general adhesion strength [36, 37]. In this paper we used a simplified version of the bioadhesion model by Decuzzi and Ferrari [38], in which the adhesion strength directly correlates with the higher chance of successful bonds of particles to the boundaries, $P$. To use this model we assume the rectangles to be oval-like particles with aspect ratio $\gamma$. Thus, for particles of circular shape,
@@
where $F^{s}$ and $T^{s}$ are coefficients that hinge upon the aspect ratio of particles $\gamma$ [38], and $\mu_{s}$ is the wall-shear stress. Eq. 12 shows that the strength of adhesion is affected by the geometric features of the particle, such as its radius $R_{equiv}$ and the aspect ratio $\gamma$.
The variation of the adhesive probability $P$ with the radius $R_{equiv}$ for three different wall shear stress regions is shown in Fig. 7(a). These regions consist of three sectors: a shear acceleration zone at the proximal end of stenosis (circles and blue solid line), a peak shear zone just before the apex of stenosis (squares and red dash-dotted line), and a shear deceleration zone at the distal end of stenosis within the recirculation area (crosses and dashed line). Zones with smaller wall shear stress show a larger adhesive probability [38]. For any of the sectors, $P$ reaches a maximum value for some intermediate optimal value of the radius and then decreases. Results not shown for rectangle particles lead to an overall increase of the adhesive probability $P$ as the aspect ratio increases.

### v2

@@ Abstract
Cardiovascular diseases are a leading cause of death globally. Among them, some are linked to stenosis, which is an abnormal narrowing of blood vessels, as well as other factors. Smart drug delivery systems based on micro- and nanoparticles are a promising method to offer non/minimal-invasive therapeutic mechanisms. Here we investigate the propensity of particles with different shapes and sizes to drift laterally (marginate) towards an occlusion area in a two-dimensional (2D) parallel plate laminar flow using the Lattice-Boltzmann method (LBM). To verify the outcomes on both sides of the stenosis, a probability of adhesion to the borders was calculated. Analysis was done on the impact of wall-shear stress on both sides of the stenosis. Our results show that rectangular particles migrate in larger amounts and earlier than circular ones.
keywords:
Lattice-Boltzmann method , Particle dynamics , Blood flow , Constricted flow
journal: Communications in Nonlinear Science and Numerical Simulation
[…]
1 Introduction
@@
Regarding the characterization of the blood, whole blood is a complex suspension made of a liquid phase, the plasma, and the cellular elements suspended within it, including red blood cells (RBCs), white blood cells, and platelets. Plasma itself behaves as a Newtonian fluid [25]; in contrast, whole blood exhibits non-Newtonian behaviour primarily due to the presence of RBCs. Nevertheless, for many applications blood can be considered a Newtonian fluid, provided that the wall shear stress observed in the regions of study are not too low [26, 27]. In such cases, the rheological differences in blood behavior are negligible and the Newtonian assumption holds true. Since this study focuses exclusively on the plasma, the use of a Newtonian model is therefore justified.
2 Numerical method
To study the motion of a particle cluster in a 2D constricted channel, the Lattice-Boltzmann method (LBM)[28, 24, 29] is adopted. The dynamics of the Newtonian fluid flow with an external force can be described by a single relaxation time via the Bhatnagar-Gross-Krook (BGK) equation [30],
@@ 2 Numerical method
Here $\rho$ and $\bf{u}$ are the fluid density and velocity, respectively; $\omega_{i}$ represents the weighting parameter of the model, where $\omega_{0}=4/9$, $\omega_{1-4}=1/9$ and $\omega_{5-8}=1/36$; ${\bf{c}}_{i}$ is the discrete velocity of the model, with $c_{0}=(0,0)$, $c_{1-4}=c(\cos[(i-1)\pi/2],\sin[(i-1)\pi/2])$, $c_{5-8}=2c(\cos[(2i-1)\pi/4],\sin[(2i-1)\pi/4])$; $c=\Delta x/\Delta t$ is the lattice speed, which was assumed to be one since $\Delta x=\Delta t=1$; $\Delta x$ is the lattice spacing; and $c_{s}=c/\sqrt{3}$ is the sound speed of the model. The last term in Eq. (1) represents the source term, which takes into account the force density driving the fluid as well as the forces exerted by particles, for which we use the Guo forcing scheme [32]. Based on this scheme, a body force density $\bf{F}$ is related to $F_{i}$ as,
@@
The local fluid density, $\rho$, and velocity, $\bf{u}$, at the lattice node are given by $\rho=\sum_{i}f_{i}$ and $\rho{\bf{u}}=\sum_{i}f_{i}{\bf{c_{i}}}+\Delta t{\bf{F}}/2$. $\bf{F}$ is the sum of the forces acting on the fluid. The LBM for a single-phase flow recovers the Navier-Stokes equation in the limit of small Knudsen number for incompressible fluids, with the fluid kinematic viscosity being $\nu=c_{s}^{2}(\tau-0.5)$ (in lattice units). The kinematic viscosity cannot be less than zero; therefore, the limitation $\tau>0.5$ must be satisfied.
The particle–fluid interaction was implemented using a force-coupling immersed boundary lattice Boltzmann method (IB-LBM), conceptually following the formulation of Safdari and Kim [33]. In this approach, the particle surface is discretized into uniformly distributed boundary nodes, whose velocities are prescribed by the rigid-body motion of the particle. At each boundary node, the surrounding fluid velocity is interpolated and compared to the particle’s velocity to evaluate the hydrodynamic forces and torque contributions. This approach differs from the dissipative-coupling method proposed by Ahlrichs and Dünweg [34], as no friction parameter is required, and from the momentum exchange method of Niu et al. [35] or the classic IBM [36], since our implementation explicitly evaluates drag and lift contributions, which are subsequently spread back to the fluid through Eq. (4), rather than via a discrete delta function.
@@
where $m_{p}$ and $I_{p}$ denote the mass and moment of inertia of the $p$ particle, ${\bf{v}}_{p}$ and $\Omega_{p}$ are the translational and rotational velocities of each particle, $\bf{F_{D}}$ represents the drag force, $\bf{F_{L}}$ is the lift force, and $\bf{T}$ the particle torque. Note that the effect of forces like virtual mass, Basset history and Magnus effect can be considered negligible. Other forces (Brownian, gravity, etc) acting on the particles are not considered in this paper. Particles’ boundary are discretized in $M$ nodes as seen in Fig. 1, whose velocities ${\bf{v}}_{b}$ are given by,
@@
with ${\bf{r}}_{b}$ and ${\bf{r}}_{p}$ being the boundary node and particle positions, respectively.
Figure 1: Representation of a single discretized circular particle (red dots) moving through the fluid lattice (white dots). Each particle boundary node is assigned a velocity $\bm{v_{b}}$. The corresponding fluid velocity at the same location, $\bm{u_{b}}$, is obtained via a four–neighbor interpolation, illustrated by the gray rectangle.
Finally, the drag and lift forces and the torque are given as,
@@
$\bf{T}$ $=$ $C_{T}\sum_{b}\left(({\bf{r}}_{b}-{\bf{r}}_{p})\times({\bf{u}}_{b}-{\bf{v}}_{p})\right)$ (10)
where $C_{D}$, $C_{L}$ and $C_{T}$ denote the drag, lift, and torque coefficients, respectively, and ${\bf{u}}_{b}={\bf{u}}({\bf{r}}_{b})$ is the fluid velocity at the boundary nodes. The summation runs over all $M$ boundary nodes. For the lift equation (9) we considered the Saffman force [37] with $C_{L}=0.5$ for the sake of simplicity, although more complex models could be used [38]. The drag coefficient is taken as a function of the particle Reynolds number, following the Schiller–Naumann correlation [39] (see also [40]). We further assume that the torque coefficient equals the drag coefficient, i.e $C_{T}=C_{D}$.
Then, the position and velocity of each particle is updated by a forward explicit Euler scheme:
${\bf{r}}_{p}(t+\Delta t)={\bf{r}}_{p}(t)+\Delta t{\bf{v}}_{p}$ (11)
${\bf{v}}_{p}(t+\Delta t)={\bf{v}}_{p}(t)+\Delta t{\bf{F}}$ (12)
with $\bf{F}=\bf{F_{D}}+\bf{F_{L}}$. The accuracy of this coupling procedure, as well as the overall implementation, has been validated against standard benchmarks, which are reported in the Supplementary Material.
@@
Fig. 2 shows the 2D rectangular domain ($H\times L=400\times 4000$ cells) considered (see A for a grid convergence study). The constriction zone or stenosis is simulated as a Gaussian function of height $h_{0}$ and variance $h_{0}^{2}$ centered at $L/2$. A Poiseuille flow with a maximum velocity $u_{max}$ at $H/2$ was considered at the inlet and an outflow boundary condition was imposed for the outlet, both achieved through the use of Zou/He type boundary conditions on the mesoscopic level [41]. The bounce-back scheme was implemented on channel top and bottom walls—including the curved Gaussian regions—to simulate the no-slip condition on the fluid, capturing the wall and inertial effects on particle-fluid motion, as discussed by Basagaoglu et al. (2008)[42].
@@
Particles had initially the same velocity as the flow at the release location. In this paper, we focus specifically on circular and rectangular rigid particles. $N$ particles were released randomly far enough from the constriction zone and at $H/4$ cells far from the lateral boundaries (gray area in Fig. 2). Care was put in making sure there is no overlapping between them. Rectangles were set initially at random orientations. For a circular particle with radius $R$ rotating around its center of mass, the angular moment of inertia is $I_{p}=m_{p}R^{2}/2$, while for a rectangular particle with sides $l_{1}$ and $l_{2}$, $l_{2}\geq l_{1}$, $I_{p}=m_{p}(l_{1}^{2}+l_{2}^{2})/12$. We consider elastic collisions between two particles, or a particle and a wall. For the collision between circular particles we suppose a hard sphere collision, whereas for the collision between two rectangles we follow Qin et al. (2020) [43]. For the last case, the collision model takes into account whether the collision involves two vertices, a vertex and an edge or two edges. In this paper, we do not consider collisions between circular and rectangular particles. To compare results between both types of particles, we define the equivalent radius for the rectangle particles as $R_{equiv}=l_{1}(1+\gamma^{2})^{1/2}/2$ where the aspect ratio is defined as $\gamma=l_{2}/l_{1}$. The fluid is initialized with a density $\rho=1$, velocity $u_{max}=0.04$, Reynolds number $Re=Hu_{max}/\nu$ and a relaxation time $\tau=0.62$, all given in lattice units. Henceforth, all results presented in this work are expressed in lattice units.
3 Results
3.1 Flow trajectories of particles
@@
The LBM was used to simulate the migration trajectories of particles as a function of their size and shape. Simulation results in Fig. 3 unveiled two distinct shape dependent particle behaviors in a constricted channel: (i) both circular and rectangular particles migrate towards the lateral boundaries; however, the rectangles do so before the circles. Additionally, (ii) rectangular particles migrate to both sides of the stenosis and in greater quantity than circles. This finding aligns with previous studies focusing on the behaviour of particles with different aspect ratios ranging from spherical to discoidal [44]. Note that both types of particles create a wake, in which others are caught and therefore experience less drag, resulting in slower motion and eventual clustering. This is less pronounced for rectangles as they displayed non-zero cumulative angular rotations favoring their dispersion, unlike the circular particles.
@@
Particle migration to the edges depends on their size and shape as their response to drag and lift forces is different. Particles are considered to have migrated to the edges if the distance from any particle node to the boundaries is less than 5 cells and they are in the intervals [L/2L/8,L/2][L/2-L/8,L/2] before and $[L/2,L/2+L/8]$ after the stenosis. Fig. 5 shows the ratio of migrated particles $N_{migr}/N$ for circles and rectangles as a function of $R_{equiv}$ before (a) and after (b) the stenosis. Note that for both occlusion percentages, $N_{migr}$ is larger for the rectangular particles than for the circular ones. Small circular particles hardly see their trajectory modified and are instead confined in the center-region of the channel. Before the stenosis, $N_{migr}$ increases with $R_{equiv}$ in all cases, while the number of circles reaching the boundaries after the stenosis decreases as their size increases. Large particles have larger residence times, Fig. 5(c), as they were less frequently trapped by the recirculating vortices. Residence time increases with the size of the occlusion zone and the recirculating vortices.
Once particles cross the stenosis, Fig. 5(b), circles are dragged by the central flow, where the velocity is at a maximum ($N_{migr}$ decreases with increasing $R_{equiv}$). For the rectangles, lift forces increase and they have a higher probability of migrating to the boundaries. Residence times do not change appreciably with $R_{equiv}$, but, in contrast to the previous case, decrease as the percentage of occlusion increases, Fig. 5(d). Very large particles can collapse/stop the flow at the stenosis by stacking [45]. It should be noted that in three-dimensional (3D) simulations, although particle accumulation may occur, the fluid velocity would not drop to zero, as there is space for the fluid to flow around the particles even if they partially obstruct the stenosis throat[46].
@@
The influence of the occlusion percentage on particle migration is shown in Fig. 6. For both types of particles with similar values of $R_{equiv}$, the number of particles that migrate to the boundaries increases with an increasing occlusion; in fact, this trend is not exclusive to artificial particles, as it was also observed by Yazdani and Karniadakis in a similar study on platelet margination in the presence of RBCs [47]. Wall-shear stress upstream of the stenosis increases with $h_{0}$, favoring migration of particles to the boundaries; conversely, post-stenosis wall-shear stress diminishes due to the presence of a recirculation zone, thus reducing the rate of particle migration, especially for circular particles, which are more prone to be dragged by the flow. The driving forces for margination are significantly altered by the constriction.
@@
Fig. 7 illustrates the influence of the number of released circular particles $N$ and the Reynolds number on the ratio of particles that migrated to the boundaries. All Reynolds numbers considered lie within the physiologically relevant range for standard arteries. Only results containing circular particles are displayed; earlier results demonstrating rectangular particles indicate that $N_{migr}$ is greater than that of circular particles. As expected, $N_{migr}$ increases with $N$ before and after the stenosis, but increasing $N$ leads to larger slow flow at the particles wake and clogging at the constriction, thus reducing the ratio of particles that migrate to the lateral boundaries with respect to the number of emitted particles. This is more evident after the stenosis. On the other hand, increasing the Reynolds number by increasing $u_{max}$ leads to a linear increase of $N_{migr}$, as the drag and lift forces are proportional to $Re$.
@@ 3.2 Probability of adhesion
There are numerous models for calculating the general adhesion strength [48, 49]. In this paper we used a simplified version of the bioadhesion model by Decuzzi and Ferrari [50], in which the adhesion strength directly correlates with the higher chance of successful bonds of particles to the boundaries, $P$. In the original model, the adhesive strength of the particle arises from the selective binding between receptors on the target surface and complementary ligands displayed on the particle surface. and the strength of adhesion is determined by geometric factors (particle volume and aspect ratio), biophysical parameters (equilibrium separation distance, maximum bonding distance, wall shear stress relative to receptor density, and the densities of ligands and receptors), and biochemical parameters (bond characteristic length and affinity constant of the ligand–receptor pair).
To use this model we assume the rectangles to be oval-like particles with aspect ratio $\gamma$, and we adopt the same threshold of 5 cells as in the margination calculation (Section 3.1), beyond which adhesion is not considered possible. Thus, for particles of circular shape,
@@
where $F^{s}$ and $T^{s}$ are coefficients that hinge upon the aspect ratio of particles $\gamma$ [50], and $\mu_{s}$ is the wall-shear stress. Eq. 14 shows that, in this formulation, adhesion strength is primarily governed by the geometric properties of the particle, namely its radius $R_{equiv}$ and the aspect ratio $\gamma$.
The variation of the adhesion probability $P$ with the radius $R_{equiv}$ for three different wall shear stress regions is shown in Fig. 8(a). These regions consist of three sectors: a shear acceleration zone at the proximal end of stenosis (circles and blue solid line), a peak shear zone just before the apex of stenosis (squares and red dash-dotted line), and a shear deceleration zone at the distal end of stenosis within the recirculation area (crosses and dashed line). Zones with smaller wall shear stress show a larger adhesive probability [50]. For any of the sectors, $P$ reaches a maximum value for some intermediate optimal value of the radius and then decreases. Results not shown for rectangular particles lead to an overall increase of the adhesive probability $P$ as the aspect ratio increases. All these observations are consistent with [50].